HOME
ESR
EXPERIMENTS
LABS etc
|
E121
E127
E132
E125
E143
laser_cooling_2021
E142
Ê128
E146
E0052
E0018
E0028
E0038
E0075
E0203_E0018
|
masses & isomeres 2025, Page 4 of 8 |
Not logged in |
 |
|
New entries since: | Thu Jan 1 01:00:00 1970 | |
Thu Apr 3 01:14:16 2025, Zac, Johan, General, Beam is back at 01:12
|
Thu Apr 3 00:49:00 2025, Zac, Johan, General, Beam lost at 00:46
|
Wed Apr 2 23:40:23 2025, JEL, ZN, General, Screenshots taken just as beam came back  
|
Wed Apr 2 23:33:43 2025, Zac, Accelerator, Beam is gone at 23:29
|
Wed Apr 2 23:27:35 2025, Zac, Johan, Accelerator, Beam is back 23:25
|
Wed Apr 2 23:15:52 2025, Zac, Johan, Accelerator, No injections from SIS
|
Wed Apr 2 22:47:25 2025, ID, HA (Radio Shahab as background), General, Lassie and Schottky TV screentshots  
|
Wed Apr 2 21:22:41 2025, ID, HA, General, Lassie Come-Home Screenshots (wuff!)  
|
Wed Apr 2 20:57:34 2025, Shahab, Detectors, 10GB cables for the DAQ rack
|
Wed Apr 2 20:15:29 2025, ID, HA, General, Lassie screenshots + Schottky TV  
|
Wed Apr 2 19:13:50 2025, ID, HA, General, Lassie screenshots and Schottky screenshot  
|
Wed Apr 2 18:31:52 2025, ID, General, FYI for morning shifts Thu
|
Wed Apr 2 18:22:30 2025, ID, General, Spill cycle distribution
|
Wed Apr 2 18:15:06 2025, ID, HA, MW, General, Lassie screenshots
|
Wed Apr 2 18:13:20 2025, ID, HA, General, RSA spectra
|
Wed Apr 2 17:46:12 2025, Ruijiu, Wolfram, Xinliang, Wangqian, Analysis, The spectrum of 194Pb82+ vs 184Pt78+  
|
Wed Apr 2 17:06:45 2025, ID, HA et al, General, Lassie screenshots
|
Wed Apr 2 16:28:14 2025, SS, ID, HA etc, DAQ, NTCAP stop/start
|
Wed Apr 2 15:46:24 2025, ID, HA,MW, General, Beam back in ESR,
|
Wed Apr 2 15:07:52 2025, Michael, Alex, Maeve, Accelerator, Still not a good beam in the ESR
|
|